### v1

# Initializing Services in Interactive ML Systems for Diverse Users

Abstract
This paper studies ML systems that interactively learn from users across multiple subpopulations with heterogeneous data distributions. The primary objective is to provide specialized services for different user groups while also predicting user preferences. Once the users select a service based on how well the service anticipated their preference, the services subsequently adapt and refine themselves based on the user data they accumulate, resulting in an iterative, alternating minimization process between users and services (learning dynamics). Employing such tailored approaches has two main challenges: (i) Unknown user preferences: Typically, data on user preferences are unavailable without interaction, and uniform data collection across a large and diverse user base can be prohibitively expensive. (ii) Suboptimal Local Solutions: The total loss (sum of loss functions across all users and all services) landscape is not convex even if the individual losses on a single service are convex, making it likely for the learning dynamics to get stuck in local minima. The final outcome of the aforementioned learning dynamics is thus strongly influenced by the initial set of services offered to users, and is not guaranteed to be close to the globally optimal outcome. In this work, we propose a randomized algorithm to adaptively select very few users to collect preference data from, while simultaneously initializing a set of services. We prove that under mild assumptions on the loss functions, the expected total loss achieved by the algorithm right after initialization is within a factor of the globally optimal total loss with complete user preference data, and this factor scales only logarithmically in the number of services. Our theory is complemented by experiments on real as well as semi-synthetic datasets.
1 Introduction
In various applications including federated learning (Li et al., 2020), crowd-sourcing (Steinhardt et al., 2016), online recommendation systems (Song et al., 2014), data about user preferences is acquired through interactions with multiple users or sources. The primary objective is to leverage this data to deliver services that cater to the diverse requirements of users. Typically, these users exhibit heterogeneity in their preferences, and recent approaches demonstrate the benefits of deploying and updating multiple, more specialized models versus using one large shared model as a one-size fits all approach; e.g., for clustered federated learning Mansour et al. (2020); Sattler et al. (2020); Ghosh et al. (2020) (Dwork et al., 2018), in the context of fair classifiers (Ustun et al., 2019), and for meta learning (Kong et al., 2020). And even when a single large model is used, it is often followed by a fine-tuning step for specialization to a group of users or tasks by retraining part of the model (Chua et al., 2021; Sun et al., 2021).
In this paper we consider ML systems that interactively learn from users across multiple subpopulations with heterogeneous data distributions. The aim is to provide a range of services, each service specialized for a group of users. These services typically work by trying to predict user preferences, and the likelihood of any user picking a service from the available candidates depends on the services’ performance for them. For example, Meta [^1] https://nickclegg.medium.com/you-and-the-algorithm-it-takes-two-to-tango-7722b19aa1c2 uses ML to predict user preferences and make personalized recommendations for its services like Facebook, Instagram, etc. Users decide whether to join recommended pages, leave likes or other forms of engagement such as scrolling activity, hours online indicating their level of satisfaction with the services. Meta would in turn plan which users each of its platforms is tailored to, and what personalized content is recommended to users—these decisions by Meta could altogether be modeled as choosing the parameters of its services.
These services are usually trained through interaction with users. Once a set of services is launched users select the service that best predicts their preferences (generally choosing the service that minimizes the particular user’s performance metric). These services are then subsequently refined based on data of current users. This sequence of (alternating) decisions by users and by services (learning dynamics) continues as more data is gathered. Broadly speaking, such a learning system could include a set of independent learners (e.g., different service providers) having an incentive to coordinate for the purpose of initializing their services, or a single learner with several deployed models (e.g., a company that owns different platforms, or a multi-model server in certain federated learning setups). This setup has been studied by Ginart et al. (2021), who discuss the effect of competition between services on user welfare, and Dean et al. (2022), who discuss the the stability of equilibria of learning dynamics in such systems.
Our goal is to design a set of services to minimize the total loss, i.e., the sum of all losses across users. However, we face the following challenges:
  • Unknown User Preferences: In practice, the presence of a service often precedes data collection. Specifically, in contexts like online recommendations, it is usually not feasible to gather user preference data without first deploying the service and observing user interactions. This means that data collection is inherently conditional on the existence of a service, challenging the conventional “data-first, model-second” paradigm.
  • Suboptimal Local Solutions: The total loss (sum of losses across all users and all services) landscape in these applications is non-convex (jointly in both sets of variables for users and services) even if the individual user losses on a single service are convex, making it likely for the learning dynamics to get stuck in local minima where the total loss can be significantly worse than the globally minimum loss. Thus, the outcomes of learning dynamics are heavily affected by initialization.
Contributions.
The following summarizes the contributions of this paper.
  • In this work, we address these challenges by designing a computationally and statistically efficient algorithm for initialization of services prior to learning dynamics. The algorithm works by adaptively selecting very few users to collect data from (via zeroth order or bandit access to their loss functions) in order to initialize the set of services.
  • We establish an approximation ratio for our algorithm: the expected total loss achieved by the algorithm right after initialization is within a factor of the globally optimal total loss in the presence of complete user preference data, and this factor scales only logarithmically in the number of services (cf. Section 3).
  • When users belong to a set of unknown demographic groups, it is desirable that the provided services do not result in unfavorable outcomes towards certain demographics (say based on gender or racial groups). A fair objective would be to minimize the maximum average loss of users across different groups. We provide an approximation ratio for our algorithm on the fair objective as well, that scales only logarithmically in the number of services (cf. Section 3).
  • In the context of linear prediction models, we study generalization properties of the initialized services to users the services have never interacted with before (cf. Section 4).
  • We empirically demonstrate the strengths of our initialization scheme through experiments on a prediction task based on 2021 US Census data, and online movie recommendation task based on Movielens10M dataset (cf. Section 5).
Notation and Terminology.
For a symmetric matrix $\mathbf{A}$ and any vector $x\in\mathbb{R}^{d}$, we denote its Mahalanobis norm by $\|x\|_{\mathbf{A}}=\sqrt{x^{\top}\mathbf{A}x}$. The generalized eigenvalues for a pair of symmetric matrices $\mathbf{A}$ and $\mathbf{B}$ are denoted by $\lambda(\mathbf{A},\mathbf{B})$, defined as the solutions of $\lambda$ for the generalized eigenvalue problems $\mathbf{A}v=\lambda\mathbf{B}v$ (Ghojogh et al., 2019; Förstner and Moonen, 2003). Specifically we use $\lambda_{\min}(\mathbf{A},\mathbf{B})$ to denote minimum generalized eigenvalue for the matrix pair $\mathbf{A},\mathbf{B}$.
The loss for a user $i$ given service $\theta\in\mathbb{R}^{d}$ is denoted by $\mathcal{L}_{i}(\theta,\phi_{i})$ where $\phi_{i}$ parameterizes the user’s preference. For a set of users $\mathcal{A}$ (e.g., $\mathcal{A}=[n]$ denotes a set of $n$ users) and a set of services $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$, the total loss is defined as $\mathcal{L}(\Theta,\mathcal{A})=\sum_{i\in\mathcal{A}}\min_{j\in[k]}\mathcal{L}_{i}(\theta_{j},\phi_{i})$.
2 Problem Setup
We consider the setting where a provider wants to design $k$ services for $n$ users with unknown diverse preferences, such that each service specializes to a subgroup of users. Each service typically uses a model parameterized by a vector $\theta\in\mathbb{R}^{d}$, which predicts users’ preferences. A user is more likely to use a service that yields higher utility by predicting their preference more accurately (corresponding to lower prediction loss).
Let the loss incurred by some user $i$ upon choosing some service parameterized by $\theta\in\mathbb{R}^{d}$ be denoted by $\mathcal{L}_{i}(\theta,\phi_{i})$. Here $\phi_{i}\in\mathbb{R}^{d}$ represents this user’s unknown preference parameter and $\mathcal{L}_{i}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is an unknown (nonnegative) function governing this user’s loss (which the provider can only evaluate at a given point; i.e., has zeroth order or bandit access to this function only, see Section 3 for more details).
For example, in the case of movie recommendations, $\phi_{i}$ could be the user’s interests across different genres of movies, $\theta$ a set of recommended movies, and $\mathcal{L}_{i}(\theta,\phi_{i})$ measures user dissatisfaction, say some function of hours spent online for this user, or a rating the user left upon using this recommendation. Users can have diverse preferences and loss geometries. The example in Figure 1 emphasizes that users not only have distinct preferences, but also experience varying degrees of dissatisfaction or satisfaction when presented with recommendations not exactly their preference.
Note that information about $\mathcal{L}_{i}$ and $\phi_{i}$ is obtained only upon interacting with the users, which we will discuss in Section 3.
Figure 1: Imagine two users who both have a strong preference for action movies (and both are content if they receive recommendations for this genre). However, User 1 also enjoys romantic comedies, but dislikes documentaries, while User 2 has the opposite preference. If User 1 is recommended a documentary, they will be dissatisfied, whereas User 2 will still remain content. Consequently, despite their shared fondness for action movies, their levels of dissatisfaction or satisfaction diverge when they encounter content from a different genre.
We make the following assumptions about the functional form of $\mathcal{L}_{i}$, and state several examples of function classes satisfying these properties.
Assumption 1 (Unique Minimizer).
The loss function satisfies the following equivalence: $\mathcal{L}_{i}(\theta,\phi_{i})=0\ \Longleftrightarrow\ \theta=\phi_{i}$.
This assumption implies that unless all users have identical preferences, there doesn’t exist a single service parameter $\theta$ that simultaneously minimizes every user’s loss. Thus providing multiple services (multiple $\theta$’s) where the users choose the one best for them is strictly better than one service for all users.
Assumption 2 (Approximate Triangle Inequalities).
For a pair of users $i,j$ there exists a finite constant $c_{ij}>0$ such that for all $\theta\in\mathbb{R}^{d}$ the following hold:
  1. (i)
    $c_{ij}\mathcal{L}_{i}(\theta,\phi_{i})\leq\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})$.
  2. (ii)
    $c_{ij}\mathcal{L}_{i}(\phi_{j},\phi_{i})\leq\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{i}(\theta,\phi_{i})$.
Here $c_{ij}$ (equal to $c_{ji}$) captures the alignment between the preferences and loss geometries of two users. Lower values of $c_{ij}$ indicate less similarity. Item $(i)$ implies that the loss for user $i$ on any service $\theta\in\mathbb{R}^{d}$ is no worse than (up to a constant factor) the sum of (a) loss of another user $j$ on using the same service, and (b) the loss of user $j$ if they were to use user $i$’s preference. The latter term (b) can be seen as measuring the similarity between the users’ preferences.
For condition $(ii)$ to hold for all $\theta\in\mathbb{R}^{d}$, it must also hold for the service that minimizes the sum of losses of both users using the same service. Suppose that users $i$ and $j$ were to exchange preferences. Then their loss would be no worse than (up to a constant factor) their minimum total loss, i.e.,
$\mathcal{L}_{i}(\phi_{j},\phi_{i})\leq\frac{1}{c_{ij}}\min_{\theta\in\mathbb{R}^{d}}\left(\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{i}(\theta,\phi_{i})\right).$
Some examples of functions (with corresponding constants) satisfying these assumptions include the following (see Appendix A for additional examples and derivations):
  • Squared error loss for linear predictors (cf. Section 4).
  • The Huber loss on the prediction error:
    $\mathcal{L}_{i}(\theta,\phi_{i})=\begin{cases}\frac{1}{2}\|\theta-\phi_{i}\|^{2},&\text{if }\|\theta-\phi_{i}\|\leq\delta\\ \delta(\|\theta-\phi_{i}\|-\frac{1}{2}\delta),&\text{otherwise.}\end{cases}$
[…]
  • The normalized cosine distance: $\mathcal{L}_{i}(\theta,\phi_{i})=1-\theta^{\top}\phi_{i}$ where $\|\theta\|_{2}=\|\phi_{i}\|_{2}=1$, with $c_{ij}=\frac{1}{2}$. This is commonly used as a similarity measure in natural language processing applications, for example finding similarity between two documents.
  • The Mahalanobis distance: $\mathcal{L}_{i}(\theta,\phi_{i})=\|\theta-\phi_{i}\|_{\Sigma_{i}}$. Different users can have different $\Sigma_{i}$ capturing their diverse loss variation (cf. Figure 1), as long as $\Sigma_{i}$ is full rank. Here $c_{ij}=\min\{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j}),\lambda_{\rm min}(\Sigma_{j},\Sigma_{i})\}$.
  • Any distance metric: This naturally follows from triangle inequality, hence $c_{ij}=1$. (Note that the examples above are also all distances, but when focusing on their special form we obtain better estimates for the $c_{ij}$.)
Suppose the users have access to $k$ services parameterized by $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$. Then, each user $i$ selects service $\theta_{l}$ that minimizes their loss, i.e. $\mathcal{L}_{i}(\Theta,\phi_{i})=\min_{l\in[k]}\mathcal{L}_{i}(\theta_{l},\phi_{i})$.
Our objective is to design $\Theta$, such that the sum of losses across users and services is minimized. We define the objective as follows:
$\mathcal{L}(\Theta,[n])=\sum_{i\in[n]}\min_{j\in[k]}\mathcal{L}_{i}(\theta_{j},\phi_{i}).$ (1)
Definition 2.1.
Define the unknown optimal set of $k$ services that minimizes the objective to be
$\Theta_{\rm OPT}:=\operatorname*{\mathrm{argmin}}_{|\Theta|=k}\mathcal{L}(\Theta,[n]).$
Specifically, $\Theta_{\rm OPT}$ defines a “clustering”, meaning a partitioning of the $n$ users into $k$ clusters. The cluster $\mathcal{B}_{m}$ is the set of all users that prefer the service $\theta_{m}$ among all the services in the optimal set $\Theta_{\rm OPT}$. In other words, $\mathcal{B}_{m}$ is defined as the set of all points such that $\mathcal{B}_{m}=\{i\in[n]\;|\;\theta_{m}=\operatorname*{arg\,min}_{\theta_{l}\in\Theta_{\rm OPT}}\mathcal{L}_{i}(\theta_{l},\phi_{i})\}$. If multiple services are equally preferred by a subpopulation, the ties are broken arbitrarily. The resulting set of clusters is denoted by $\mathcal{C}(\Theta_{\rm OPT})=\{\mathcal{B}_{1},\ldots,\mathcal{B}_{k}\}$.
The are several statistical and computational challenges to this problem.
Challenge 1.
Since preferences $\{\phi_{i}\}_{i\in[n]}$ and loss functions $\{\mathcal{L}_{i}\}_{i\in[n]}$ are unknown, and the provider only has zeroth order or bandit feedback access, estimating the objective function $\mathcal{L}(\Theta,[n])$ usually needs a lot of data collected uniformly across the users. This large amount of data is needed before services can be deployed, yet as stated earlier, we are in the situation where we have no data until we deploy services and observe user interactions. Our limited access to user information (via limited queries) make our setting challenging.
Challenge 2.
In addition to the first challenge, the loss function is non-convex and iterative minimization approaches from a random initialization are susceptible to getting stuck in arbitrarily poor local optima. Alternatively, computing the optimal clustering first and then finding the best service for each cluster would be NP-Hard.
Despite these challenges, in Section 3, we propose an algorithm that is both statistically and computationally efficient, and admits an approximation ratio with respect to the globally optimal value, i.e., $\mathcal{L}(\Theta_{\rm OPT},[n])$.
Algorithm 1 AcQUIre- Adaptively Querying Users for Initialization
1:  Input: Set of users $[n]$, number of services $k$.
2:  Choose a user $i$ uniformly randomly from $[n]$.
3:  Query user $i$’s preference $\phi_{i}$, set the first service $\Theta_{1}=\phi_{i}$.
4:  for $t\in\{2,\ldots,k\}$ do
5:     User behavior: Collect user losses on existing services $\Theta_{t-1}$ : $\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[n]}$.
6:     User Selection: Sample $l$ from $[n]$ with probability $P(l=i)\propto\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})$.
7:     New service: Query user $l$’s preference $\phi_{l}$.
8:     $\Theta_{t}=\Theta_{t-1}\cup\phi_{l}$.
9:  end for
10:  Return $\Theta_{k}$
3 Algorithm & Main Results
In this section, we present our initialization algorithm with guarantees, Algorithm 1, and describe how the steps of the algorithm arise naturally in the interactive systems under consideration. Since collecting preference data uniformly across all the $n$ users can be prohibitively expensive, our goal is to get data from a minimal number of users.
Each iteration of the loop in the algorithm adds a service sequentially and the loop terminates when there are $k$ services, where $k$ is a predetermined parameter for the algorithm. We focus on the loop (lines 4-8) in Algorithm 1.
Suppose at time $t-1$, the set $\Theta_{t-1}$ is the set of current $t-1$ services. Then, at time step $t$ the following steps take place.
  • User behavior (line 5). Given the list of services $\Theta_{t-1}$ users are assumed to choose the best service that minimizes user risk. Users report their losses with respect to the service they choose from the set of existing services $\{\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[n]}$. In practice, this step requires deploying the services and collecting signals of engagement and utility to determine the loss associated with each user, under the behavioral assumption that users are rational agents that choose the best available service. The provider thus needs to measure each user’s loss only in their single chosen service.
  • User selection (line 6) A new user $l$ is selected with probability proportional to $\mathcal{L}_{i}(\Theta_{t-1},\phi_{l})$. This ensures that users that are currently poorly served by existing services are more likely to be selected.
  • New Service (line 7-8). Given a selected user $l$, the algorithm queries the preference $\phi_{l}$ of the user and centers the new service at that preference $\theta_{t}=\phi_{l}$. In practice, this step requires acquiring data about the user in order to learn their preference (this is needed for only $k$ total users throughout the algorithm). For example, data may be acquired by incentivizing the selected users, via offering discount coupons or free premium subscriptions (Hirnschall et al., 2018).
With each iteration the loss of each user is non-increasing; the previous services remained fixed and a user would switch to a new service only if it improves quality or equivalently decreases loss. Since at each iteration, a new service is added, the process terminates after $k$ steps. Since it is costly to offer and maintain too many different services, we typically have $k\ll n$.
We now discuss the theoretical properties of the set of services we get at the termination of Algorithm 1.
Theorem 1.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and unknown associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 1 and 2. Let $\Theta_{\rm OPT}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the total loss and $\mathcal{C}(\Theta_{\rm OPT})$ the resulting partitioning of users (Definition 2.1). If Algorithm 1 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
$\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[n])]$ $\leq K_{\rm OPT}(2+\log k)\cdot\mathcal{L}(\Theta_{\rm OPT},[n]),$
where the expectation is taken over the randomization of the algorithm and $K_{\rm OPT}$ is equal to
$\max_{\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})}\frac{4}{\min\limits_{j\in\mathcal{B}}\sum\limits_{i\in\mathcal{B}}c_{ij}}\left(\max_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right).$ (2)
A detailed proof is presented in Appendix B; we summarize the main ideas here. The intuition is that a chosen user’s preference is typically a good representative for other users in its cluster. Thus adding a service parameterized by the chosen user’s preferences generally reduces the losses of users in this cluster. Subsequently we are less likely to pick another user from the same cluster. The $\log k$ factor is due to clusters from which users were never picked.
A similar proof approach was used by Arthur and Vassilvitskii (2007) in the context of the $k$-means problem, by sequentially placing centers on known points sampled with probability proportional to the point’s squared distance to its closest existing center. A key novelty of our analysis is to capture the alignment of diverse loss geometries across users in a large class of functions, specifically understanding how user similarities $c_{ij}$ affect the approximation ratio. [^2] In addition, as stated earlier, we tackle the lack of prior information on $\phi_{i}$ and the function form of $\mathcal{L}_{i}$—this challenge is particular to our setting and does not arise in related standard problems of clustering ($k$-means, $k$-mediods) or resource allocation (facility location problems).
Key characteristics of $K_{\rm OPT}$:
The following are essential characteristics of the term $K_{\rm OPT}$.
  1. (i)
    All terms in $K_{\rm OPT}$ depend on the local clusters in the unknown optimal clustering $\mathcal{C}(\Theta_{\rm OPT})$.
  2. (ii)
    The constant $\min_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}c_{ij}$ captures the user whose loss geometry is least similar to the average loss geometry of the cluster they belong to (recall Assumption 2.i).
  3. (iii)
    The constant $\max_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}$ captures the user whose preference is least similar to the optimal service parameter of the cluster they belong to (recall Assumption 2.ii).
  4. (iv)
    Even within a cluster all terms are averages, so a few poorly aligned pairs of users don’t hurt the bound if the cluster sizes are large.
Fair objective.
While minimizing the total loss is beneficial from the provider’s point of view in keeping users satisfied on average, it is undesirable in human-centric applications if the provided services result in unfavorable or harmful outcomes towards some demographic groups.
Suppose the $n$ users come from $m$ unknown different demographic groups ($m$ is typically small, say racial groups, gender). We denote the groups as $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. The fairness objective is defined as the maximum average loss suffered by any group:
$\Phi(\Theta,\mathcal{A})=\max_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}.$ (3)
Ghadiri et al. (2021) defined this objective in the context of fair $k$-means where the points and group identities are known and gave a non-constructive proof that if a $c-$approximate solution for $k$-means exists, it is $m\cdot c$-approximate for fair $k$-means. In our setup, where both user preferences and group identities are unknown, the set of services initialized by Algorithm 1 enjoys guarantees on optimizing the fairness objective as stated below.
Theorem 2.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and unknown associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 1 and 2. Suppose these users belong to $m$ unknown demographic groups $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. Let $\Theta_{\rm fair}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the fairness objective $\Phi$ given in (3). If Algorithm 1 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
$\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq m\cdot K_{\rm OPT}\cdot\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}\cdot(2+\log k)\cdot\Phi(\Theta_{\rm fair},\mathcal{A}),$
where the expectation is taken over the randomization of the algorithm and $K_{\rm OPT}$ is defined in (2).
If the group identities are known a priori to the provider, we can modify Algorithm 1 to get rid of the factor $\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}$ in the approximation ratio, as discussed in Appendix C.
4 Generalization in Linear Predictors
In practical settings, a provider would want to design services that not only keep the subscribed users satisfied but also attract new users to subscribe by generalizing the services to users it has never interacted with before. Now instead of considering $n$ users, suppose that each $i\in[N]$ represents a subpopulation with its own (sub-Gaussian) distribution of features, and the provider can interact with finite samples $n_{i}$ from these distributions. A question that arises is whether we can deal with this finite-sample-from-subpopulations scenario, and how does the number of samples affect the algorithm’s output to unseen users. In this section, we answer this question for the special case of linear predictors (i.e., regression loss).
In this section we restrict ourselves to the special case of linear prediction tasks, where the goal is to accurately predict the score of a user as a linear function of their features. The score for a user in the $i^{\rm th}$ subpopulation with zero-mean random feature $x\in\mathbb{R}^{d}$ is generated as $y=\phi_{i}^{\top}x$ where both the true linear regressor $\phi_{i}\in\mathbb{R}^{d}$ and the feature covariance $\mathbb{E}_{x}[xx^{\top}]=\Sigma_{i}$ are unknown. Suppose a service uses a linear regressor $\theta\in\mathbb{R}^{d}$, to predict the score for this user as $\theta^{\top}x\in\mathbb{R}$. The loss for this subpopulation for this service is defined as the expected squared error between the predicted and actual scores, i.e., $\mathcal{L}_{i}(\theta,\phi_{i})=\mathbb{E}_{(x,y)}[(\theta^{\top}x-y)^{2}]=\|\theta-\phi_{i}\|^{2}_{\Sigma_{i}}$.
Assumption 3.
For subpopulation $i$, features are independent draws from a zero-mean sub-Gaussian distribution. For a random feature $x\in\mathbb{R}^{d}$ and for any $u\in\mathbb{R}^{d}$, such that $\|u\|_{\Sigma_{i}}=1$, $u^{\top}x\in\mathbb{R}$ is sub-Gaussian with variance proxy $\sigma_{i}^{2}$. [^3] A random variable $x$ is sub-Gaussian with variance proxy $\sigma^{2}$ if $\mathbb{E}[\exp(\lambda x)]\leq\exp(\frac{\sigma^{2}\lambda^{2}}{2})\;\forall\lambda\in\mathbb{R}$.
Assumption 4.
We assume that the decision to choose between different services happens at a subpopulation level and not an individual level.
To illustrate Assumption 4, consider the example of a provider that offers personalized services to schools (subpopulations) such as online library resources wherein the service provider queries students about the experience. Each school typically has a considerable number of students, but only a subset of them may actively respond to such queries. Once a school selects the service, it is made available to all students, and the provider could implement a system where students are encouraged to fill out a feedback form after using their service.
Suppose only $n_{i}$ users from subpopulation $i$ are subscribed to the services. Thus, upon choosing a service parameterized by $\theta\in\mathbb{R}^{d}$ the provider observes an empirical loss, which is given by
^$\widehat{\mathcal{L}}_{i}(\theta,\phi_{i})=\frac{1}{n_{i}}\sum_{j\in[n_{i}]}(\theta^{\top}x_{i}^{j}-y_{i}^{j})^{2},$
where $\{(x_{i}^{j},y_{i}^{j})\}_{j\in[n_{i}]}$ are private unknown features and scores of the users. We stress that the service gets to see the value of the user loss function at the deployed $\theta$ (bandit feedback), but not the features of each subpopulation.
Assumption 5.
The number of users from each subpopulation is greater than the dimension of the linear predictor, i.e. $n_{i}\geq d$ for all $i\in[N]$.
In this setting, given a set of services parameterized by $\Theta_{t-1}=\{\theta_{1},\ldots,\theta_{t-1}\}$ the Steps 5-6 of Algorithm 1 proceed with these finite sample averages ^$\widehat{\mathcal{L}}_{i}(\Theta_{t-1},\phi_{i})=\min_{j\in[t-1]}\widehat{\mathcal{L}}_{i}(\theta_{j},\phi_{i})$. In Step 7, multiple ways can be adopted by the provider to estimate $\phi_{i}$. Users can be given incentives to provide a batch of feature score pairs, or gradient free methods can be used to estimate the optimal solution to the regression problem. The generalization guarantee of Algorithm 1 to the total expected loss is stated below.
Theorem 3.
Suppose users belong to $N$ subpopulations satisfying Assumptions 3, 4, and 5. Let $\{n_{i}\}_{i\in[N]}$ denote the number of samples per subpopulation. Let $\Theta_{k}$ be the output of Algorithm 1 using finite sample estimates ^$\widehat{\mathcal{L}}_{i}(\cdot,\phi_{i})$ and $\Theta_{\rm OPT}$ be the optimal solution of the expected loss. Then, for any $\epsilon\in(0,1)$, if $n_{i}=\Omega(\frac{\sigma_{i}^{4}\sqrt{N}\log{(2/\delta)}}{\epsilon^{2}})$ for all $i\in[N]$, the following inequality holds with probability at least $1-\delta$:
$\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[N])]\leq\frac{1+\epsilon}{1-\epsilon}K_{\rm OPT}(2+\log k)\mathcal{L}(\Theta_{\rm OPT},[N]),$
where $K_{\rm OPT}$ is as defined in (2) and
$c_{ij}=\frac{1}{2}\min\ \left\{\lambda_{\mathrm{min}}(\widehat{\Sigma}_{i},\widehat{\Sigma_{j}}),\frac{1}{\lambda_{\mathrm{min}}(\widehat{\Sigma}_{i},\widehat{\Sigma_{j}})}\right\},$
with
$\widehat{\Sigma}_{i}=\frac{1}{n_{i}}\sum_{l\in[n_{i}]}x_{i}^{l}(x_{i}^{l})^{\top}\quad\text{and}\quad\widehat{\Sigma}_{j}=\frac{1}{n_{j}}\sum_{l\in[n_{j}]}x_{j}^{l}(x_{j}^{l})^{\top}.$
The proof is presented in Appendix D; we provide a brief overview here. We apply the Chernoff bound to the difference between the empirical loss and expected loss. Note that $(i)$ even if the same set of services are provided, the loss minimizing service for the empirical loss may be different from the expected loss for any subpopulation, and $(ii)$ the optimal set of services for the total empirical loss and the total expected loss are different. Handling these carefully, and utilizing Theorem 1 concludes the proof.
Remark 1.
Note that $\widehat{\Sigma}_{i}$ is the empirical feature covariance of subpopulation $i$. The term $c_{ij}$ captures the alignment between two subpopulation’s loss geometry, and here is equal to half of the minimum generalized eigenvalue of the empirical feature covariances of the the respective subpopulations. This quantity is the largest constant satisfying Assumption 2 (cf. Appendix D).
5 Experiments
We empirically demonstrate [^4] All our code is available at https://anonymous.4open.science/r/InteractiveSystemInitialization-2A33 the benefits of our algorithm on a commute time prediction task based on 2021 US Census data [^5] https://www.census.gov/programs-surveys/acs/data.html. and a semi-synthetic movie recommendation task on the MovieLens10M dataset. In both the tasks, our algorithm starts without access to any data and our goal is to evaluate how well the sequential acquisition of data, subsequent learning and deployment of learnt models performs on the whole dataset.
Census Data.
We consider the task of predicting daily work commute times, based on 2021 US census data from folktables (Ding et al., 2021). We illustrate a potential use case: the provider is a transport authority offering services in the form of personalized podcasts. If the duration of service is similar to the commute time of a user, that user will be able to consume the media while travelling to work. Hence an accurate prediction of the commute time may be useful in providing services tailored to the users.
The data set has $n=7025$ subpopulations defined by the intersection of gender, race and zip code. We use $k$ linear predictors (as services) with private user features (education, income, transportation mode, age etc). Refer to Appendix E for a detailed description of data pre-processing. As described in Section 4 , the user loss is squared prediction error of their commute time. The features and commute times are unknown a priori to the the provider. At time step $t\leq k$, suppose there exists a set of services $\Theta_{t-1}$. The provider observes the losses (squared prediction error as discussed in Section 4) across different subpopulations {^$\{\widehat{\mathcal{L}}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[N]}$. Based on these observed losses, a subpopulation $l$ is sampled with probability ^$\propto\widehat{\mathcal{L}}_{l}(\Theta_{t-1},\phi_{l})$. Upon selecting a subpopulation the provider surveys the population and acquires relevant data for it. Subsequently the provider uses the acquired data to estimate the sub-population preference via ordinary least squares: $\widehat{\phi}_{l}=\min_{\theta\in\mathbb{R}^{d}}\widehat{\mathcal{L}}_{l}(\theta,\phi_{l})$. The list of services is updated $\Theta_{t}=\Theta_{t-1}\cup\widehat{\phi}_{l}$ and the process is repeated until $k$ services are initialized.
Figure 2: Average loss of different demographics with services initialized by Algorithm 1. Note that Algorithm 1 has much lower losses as compared to the baseline for the each of the demographic groups.
Evaluation.
We compare our algorithm against the baseline where $k$ subpopulations are uniformly selected at random and the services initialized as their learnt regression parameters. The subpopulations belong to the following demographic groups [^6] These group labels are taken from the publicly available 2021 US Census data. (i) Based on sex: male or female, (ii) Based on race: White, Black, American Indian , Alaskan Native, Asian, Pacific Islander, Others. As shown in Figure 2, Algorithm 1 reduces the average loss for every demographic group. We also observe a 15-25% and 7-25% improvement on the fair objective—which is defined in (3)—for sex and racial group demographics, respectively, as the number of services varies (see Figure 4 in Appendix E for detailed plots). We emphasize that this result is observed despite the algorithm not knowing the identity of the demographic groups whilst collecting data.
Figure 3: Total excess error $\sum_{i\in[N]}\mathcal{L}_{i}(V,V_{i})$, where $\mathcal{L}_{i}$ is defined in (4), versus the number of services $d$ (i.e., the number of recommendation models) for Algorithm 1 versus the uniform initialization baseline.
Movie Recommendations.
We conduct a semi synthetic experiment based on the widely used Movielens10M data set (Harper and Konstan, 2015) containing 10000054 ratings across 10681 movies by 71567 viewers. We pre-process the data set (cf. Appendix E) and divide viewers into $N=1000$ subpopulations and retain only the top $m=200$ most watched movies. We compare the performance of Algorithm 1—which initializes a set of recommendation models where user preferences are unknown a priori—to a uniform initialization baseline. For subpopulation $i\in[N]$, let the solution to the following optimization problem denote the unknown user and item embeddings:
$(U_{i},V_{i})=\operatorname*{arg\,min}_{(U,V)\in\mathbb{R}^{n_{i}\times d}\times\mathbb{R}^{d\times m}}\|(UV)_{\Omega_{i}}-r_{i}\|_{2}^{2},$
where $r_{i}$ is the true user ratings and $\Omega_{i}$ is the list of movies rated by the user for subpopulation $i$. User dissatisfaction for a recommendation model, parameterized by $V\in\mathbb{R}^{d\times m}$, is captured by the excess error, namely
$\mathcal{L}_{i}(V,V_{i})=\|(U_{i}V)_{\Omega_{i}}-r_{i}\|_{2}^{2}-\|(U_{i}V_{i})_{\Omega_{i}}-r_{i}\|_{2}^{2}.$ (4)
We chose $d=5$ for the experiment. As described for the census experiment, Algorithm 1 iteratively selects subpopulations with probability proportional to their current excess error, and then solves the matrix factorization problem for the selected subpopulation, and adds the learnt item embedding to the list of services.
As shown in Figure 3, the total excess error achieved by Algorithm 1 is significantly lower than uniformly sampling users for their preferences. It is interesting to note that to achieve the same total excess error for five services initialized by our algorithm, uniform data collection and initialization would need 30 services.
6 Related Work
Retention.
User retention in machine learning systems is closely related to the decision dynamics between the provider and users studied in the single service setting by Hashimoto et al. (2018) and Zhang et al. (2019), and multiple services setting by Dean et al. (2022). In settings with multiple sub-populations of users of different types, the question of retention has been explored in parallel with the issue of fairness. The final outcome of such dynamics are heavily influenced by the initial data available to provider and the initial configuration (or partitioning) of the users across the offered services. Our work addresses these issue with theoretical guarantees.
Multiple Model Specialization.
Our work also considers the inherent heterogeneity in user preferences, especially prominent in distributed learning settings where data sources are users’ personal devices. For applications such as recommendation systems and personalized advertising, the utilization of multiple specialized models can yield improved predictions and outcomes. Users are often grouped into clusters, each representing unique interests and preferences. For instance, in movie recommendation systems, these clusters could represent users interested in different movie genres, or different combination of features. This approach has been adopted in recent federated learning formulations such as clustered federated learning (Sattler et al., 2020; Mansour et al., 2020) and is integral to the federated learning model of Ghosh et al. (2020), where users choose models for which to provide updates. This model specialization is also explored in the mixture of experts literature (see, e.g., Masoudnia and Ebrahimpour (2014)), which uses multiple ‘expert’ models to enhance accuracy and robustness, assigning inputs to the most suitable expert based on their features through a gating mechanism.
Clustering.
The $k$-means clustering problem is one of the most commonly encountered unsupervised learning problems. Given a set of known $n$ points in Euclidean space, the goal is to partition them into $k$ clusters (each characterized by a center), such that the sum of square of distances to their closest center is minimized. Dasgupta (2008) and Aloise et al. (2009) showed that the $k$-means problem is NP-Hard.
The most popular heuristic for $k$-means is Lloyd’s algorithm (Lloyd, 1982), often referred to as the "$k$-means algorithm", which proceeds by randomly initializing $k$ centers and then uses iterative updates to find a locally optimal $k$-means clustering, which can be arbitrarily bad compared to the globally optimal clustering. The performance of the $k$-means algorithm relies crucially on the initialization. Arthur and Vassilvitskii (2007) and Ostrovsky et al. (2013) proposed an elegant polynomial time algorithm for initializing centers, known as $k$-means++. Arthur and Vassilvitskii (2007) proved that the expected cost of the initial clustering obtained by $k$-means++ is at most $8(2+\log k)$ times the cost of optimal $k$-means clustering and showed a family of $k$-means instances where the approximation ratio is $2\log k$ showing their analysis is almost tight. Makarychev et al. (2020) improved the analysis to show an approximation ratio $5(2+\log k)$ and showed a family of instances with $5\log k$ approximation ratio, thus showing $k$-means++ is tight.
Our work generalizes the analysis of Arthur and Vassilvitskii (2007) to the setting where user’s preferences are represented as unknown points and the loss functions are unknown and not necessarily identical as long they satisfy Assumptions 1 and 2. Recent works in clustering (Kanungo et al., 2002; Ahmadian et al., 2019; Lattanzi and Sohler, 2019) provide a constant approximation ratio, and although they are polynomial time algorithms, these methods are data inefficient and rely on knowing the points a priori. Finally, another common clustering algorithm is Gaussian mixture models (GMMs), where each cluster is represented by a Gaussian distribution, thus clusters themselves can be viewed as ellipsoidal (which is a different goal from ours even we choose Mahalanobis distances as losses, for us in this case every point (user) is associated with a different loss geometry).
Facility Location Problem.
Our algorithm has some resemblances with the facility location problem where there are $n$ users, and a provider can set up at most $k$ facilities at one of $m$ candidate locations. This is also often known as the $k$-medoids objective (Arora et al., 2016). One of the key differences is the provider can choose from $m$ predecided locations, compared to our setting where the optimization space is infinite. However our algorithm initializes only at $k$ of the $n$ user preferences, hence it can be viewed as the candidate locations simply being the user preferences and thus $m=n$. A typical greedy algorithm for the $k$-medoids objective proceeds by evaluating the marginal decrement in total loss for all possible candidates and selects the candidate with maximal loss reduction. Thus (i) the algorithm would need to know all $\{\phi_{1},\ldots,\phi_{n}\}$ apriori, and (ii) deploy $n$ services to obtain all $n^{2}$ function evaluations $\mathcal{L}_{j}(\phi_{i},\phi_{j}),\;\forall i,j\in[n]$. This is infeasible in applications like online recommendation systems where $n$ is very large and typically a provider has the capacity to deploy only $k\ll n$ services. Hence a $k$-medoids like objective will not be reasonable in our setup with incomplete information.
Bose et al. (2022) studied the case where the provider had prior access to only $N<n$ users’ utility functions before deploying services and the goal was to minimize the worst case total (over all $n$ users) loss. However their results assume they can solve a computationally hard mixed integer program optimally.
Preference Learning.
Given a fixed set of items, Tatli et al. (2022) focus on learning preferences distributions of users. Canal et al. (2022) extend this to the setting where the user losses are given by an identical metric, and they learn both preferences and the metric efficiently. Our setting focuses on the design of services rather than learning preferences over a fixed set of services, and we also allow each user to potentially have a different loss geometry; e.g., Mahalanobis distances with different matrices (where the sublevel sets of the loss are ellipsoids with different principal axes and orientations).
7 Discussion
It would be interesting to study how strategic user behavior affects the data acquisition and services’ initialization process. For example, users may report losses that are higher than their true losses in the hopes of being selected by the provider. This in turn can lead to a biasing of the service towards these particular untruthful users. It an interesting line of questioning centers on the design of algorithm modifications or the design explicit incentives to help avoid this, for example, by encouraging truthful reporting and also voluntary participation in the case that the designed mechanisms (algorithmic or incentive based) lead to users ‘opting out’ of all services.
On a related note, seemingly benign non-strategic behavior can also lead to degradation in performance. Specifically, there are many open questions relating to quantifying the robustness of the proposed initialization algorithm to noisy observations, perturbations, or outliers in the finite sample case when the feature distribution is heavy-tailed.
Finally, it is interesting to study the improvement in the total loss not only after initialization as was done in this paper, but also after running iterative learning dynamics, alternating service updates and users’ choice updates. It is known that the total loss will not increase and the iterations converge to a local minimum of the total loss, but further quantifying this improvement, as well as other properties of the iterative updates (including whether they can be carried out in our low-information setting) is left for future work.
Acknowledgements
Dean, Fazel, Morgenstern and Ratliff were supported by NSF CCF-AF 2312774/2312775. Additionally, Morgenstern’s work was supported by NSF CCF 2045402. Dean’s research was supported by a gift from Wayfair, a LinkedIn research award, and NSF OAC 2311521. Ratliff’s research was supported by NSF CNS 1844729 and Office of Naval Research YIP Award N000142012571. Fazel was supported in part by awards NSF TRIPODS II 2023166, CCF 2007036, CCF 2212261.
Appendix A Mappings satisfying Assumptions 1 and 2
In this appendix, we provide more details on examples that satisfy Assumptions 1 and 2. For the squared error loss for linear predictors, we refer the reader to Lemma 5 for a derivation of $c_{ij}$. The remainder of the example classes are detailed below.
A.1 Huber Loss
Recall that the Huber loss is defined by
$\mathcal{L}_{i}(\theta,\phi_{i})=\begin{cases}\frac{1}{2}\|\theta-\phi_{i}\|^{2}&\text{if }\|\theta-\phi_{i}\|\leq\delta,\\ \delta(\|\theta-\phi_{i}\|-\frac{1}{2}\delta)&\text{otherwise}.\end{cases}$
Note that the Huber loss varies quadratically when the error $\|\theta-\phi_{i}\|\leq\delta$ and linearly otherwise. We will refer to these as quadratic and linear regime of the Huber loss subsequently.
We divide the derivation into 2 subcases.
Case 1. $\|\theta-\phi_{i}\|\leq 2\delta$. Within Case 1, there are several subcases to consider. We analyze each one separately.
  • At least one of the terms on the right is linear: Under the condition of Case 1, $\mathcal{L}_{i}(\theta,\phi_{i})\leq\frac{3}{2}\delta^{2}$. Note that the Huber loss is monotonicly increasing in the prediction error. The function value in the linear in $\|\theta-\phi_{i}\|$ is at least $\frac{1}{2}\delta^{2}$. Thus, if one of the terms in $\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})$ were in the linear regime, the $\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})\geq\frac{1}{2}\delta^{2}$ and $\frac{1}{3}\mathcal{L}_{i}(\theta,\phi_{i})\leq\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})$.
  • Both terms on the right are quadratic: In this sub-case, it needs to be shown when both terms of $\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})$ are in the quadratic regime the bound still holds.
    By the triangle inequality on norms, we have that
    $\|\theta-\phi_{i}\|\leq\|\theta-\phi_{j}\|+\|\phi_{i}-\phi_{j}\|.$
    Using the power mean inequality, namely
    $a\leq b+c\implies a^{2}\leq 2(b^{2}+c^{2}),$
    on the triangle inequality, we deduce the following implication:
    $\|\theta-\phi_{i}\|\leq\|\theta-\phi_{j}\|+\|\phi_{i}-\phi_{j}\|\ \implies\ \frac{1}{2}\|\theta-\phi_{i}\|^{2}\leq\|\theta-\phi_{j}\|^{2}+\|\phi_{i}-\phi_{j}\|^{2}.$
    If the $\mathcal{L}_{i}(\theta,\phi_{i})$ is quadratic, we have that $\frac{1}{2}\mathcal{L}_{i}(\theta,\phi_{i})\leq\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})$. Now, let $\mathcal{L}_{i}(\theta,\phi_{i})$ be linear. Then, we deduce that
    $\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j})=\frac{1}{2}(\|\theta-\phi_{j}\|^{2}+\|\phi_{i}-\phi_{j}\|^{2})\geq\frac{1}{4}\|\theta-\phi_{i}\|^{2}.$
    Since the equation $\frac{1}{4}x^{2}-\frac{1}{3}x+\frac{1}{6}=0$ has no real roots, the expression is always positive. Hence, by substituting $x=\frac{\|\theta-\phi_{i}\|}{\delta}$, we have that
    $\frac{1}{4}\|\theta-\phi_{i}\|^{2}\geq\frac{1}{3}\delta(\|\theta-\phi_{i}\|-\frac{1}{2}\delta)=\frac{1}{3}\mathcal{L}_{i}(\theta,\phi_{i}).$
    Therefore, we deduce that
    $\frac{1}{3}\mathcal{L}_{i}(\theta,\phi_{i})\leq\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j}).$
Case 2. $\|\theta-\phi_{i}\|>2\delta$. First, observe that
$\frac{1}{3}\mathcal{L}_{i}(\theta,\phi_{i})$ $=\frac{1}{3}\delta(\|\theta-\phi_{i}\|-\frac{1}{2}\delta)$
$=\frac{1}{3}\delta\|\theta-\phi_{i}\|+\frac{1}{6}\delta\|\theta-\phi_{i}\|-\frac{1}{6}\delta\|\theta-\phi_{i}\|-\frac{1}{6}\delta^{2}$
$\leq\frac{1}{3}\delta\|\theta-\phi_{i}\|+\frac{1}{6}\delta\|\theta-\phi_{i}\|-\frac{1}{3}\delta^{2}-\frac{1}{6}\delta^{2}$
$\leq\frac{1}{2}\delta\|\theta-\phi_{i}\|-\frac{1}{2}\delta^{2},$
where the second to last inequality follows from the fact that $\|\theta-\phi_{i}\|>2\delta$. By the triangle inequality on norms, we have that
$\|\theta-\phi_{i}\|\leq\|\theta-\phi_{j}\|+\|\phi_{i}-\phi_{j}\|.$
In turn, this implies either $\|\theta-\phi_{j}\|$ or $\|\phi_{i}-\phi_{j}\|$ is greater than $\frac{1}{2}\|\theta-\phi_{i}\|$. Without loss of generality, assume $\|\theta-\phi_{j}\|\geq\frac{1}{2}\|\theta-\phi_{i}\|\geq\delta$. Thus
$L_{j}(\theta,\phi_{j})=\delta(\|\theta-\phi_{j}\|-\frac{1}{2}\delta)\geq\frac{1}{2}\delta\|\theta-\phi_{i}\|-\frac{1}{2}\delta^{2}\geq\frac{1}{3}\mathcal{L}_{i}(\theta,\phi_{i}).$
Since $\mathcal{L}_{j}(\phi_{i},\phi_{j})\geq 0$, we have that
$\frac{1}{3}\mathcal{L}_{i}(\theta,\phi_{i})\leq\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{j}(\phi_{i},\phi_{j}).$
A.2 The normalized cosine distance
Recall that the normalized cosine distance is given by
$\mathcal{L}_{i}(\theta,\phi_{i})=1-\theta^{\top}\phi_{i}\quad\text{where}\quad\|\theta\|_{2}=\|\phi_{i}\|_{2}=1.$
Therefore, we have that
$1-\theta^{\top}\phi_{i}=\frac{1}{2}(\|\theta\|_{2}^{2}+\|\phi_{i}\|_{2}^{2}-2\theta^{\top}\phi_{i})=\frac{1}{2}\|\theta-\phi\|_{2}^{2}.$
Assumption 2 is satisfied with $c_{ij}=\frac{1}{2}$ by using triangle inequality followed by power mean inequality.
A.3 Mahalanobis distance
Consider the Mahalanobis distance which is defined by
$\mathcal{L}_{i}(\theta,\phi_{i})=\|\theta-\phi_{i}\|_{\Sigma_{i}},$
where $\Sigma_{i}$ is full rank and $c_{ij}=\min\{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j}),\frac{1}{\lambda_{\rm min}(\Sigma_{i},\Sigma_{j})}\}$. The derivation is similar to proof of Lemma 5.
Appendix B Proof of Theorem 1
We note that the parameters of the services initialized by Algorithm 1 are a subset of the users’ unknown preferences. This allows us to define the notion of covering.
Definition B.1.
Let $\Theta\subset\{\phi_{1},\ldots,\phi_{n}\}$ be a set of services. A cluster $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$ is said to be covered by $\Theta$ if there exists $i\in\mathcal{B}$ such that $\phi_{i}\in\Theta$. If no such $i\in\mathcal{B}$ exists, then the cluster $\mathcal{B}$ is said to be uncovered.
The proof idea from here on is to show that there exists an approximation ratio $K_{\rm OPT}$ for the covered clusters, and is shown in Lemma 1 and 2.
Lemma 1.
Let $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$ be an arbitrary cluster in $\mathcal{C}(\Theta_{\rm OPT})$, and let $\theta\in\mathbb{R}^{d}$ be a service centered on the preference of a user $j$ chosen uniformly at random from $[n]$. The expected loss of users in $\mathcal{B}$, conditioned on $j\in\mathcal{B}$, satisfies
$\mathbb{E}_{\theta}[\mathcal{L}(\{\theta=\phi_{j}\},\mathcal{B})\;|\;j\in\mathcal{B}]\leq\left(\max_{j\in\mathcal{B}}\frac{2}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}\right)\mathcal{L}(\Theta_{\rm OPT},\mathcal{B}).$
Proof.
Since we choose a user from $\mathcal{B}$, the conditional probability that we choose some fixed $\phi_{j}$ as the parameter for the service is precisely $P(\theta=\phi_{j}\;|\;\theta\in\mathcal{B})=\frac{1}{|\mathcal{B}|}$. Let $\mathcal{J}(\mathcal{B})$ denote the best service covering all the points in $\mathcal{B}$, and then compute
$\mathbb{E}_{\theta}[\mathcal{L}(\{\theta=\phi_{j}\},\mathcal{B})\;|\;j\in\mathcal{B}]=\sum_{j\in\mathcal{B}}P(\theta=\phi_{j})\mathcal{L}(\theta,\mathcal{B})=\sum_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\phi_{j},\phi_{i}).$
Using Assumption 2.(ii) with $\theta=\mathcal{J}(\mathcal{B})$, we have that
$\mathbb{E}_{\theta}[\mathcal{L}(\{\theta=\phi_{j}\},\mathcal{B})\;|\;j\in\mathcal{B}]$ $\leq\sum_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\left(\frac{1}{c_{ij}}\left(\mathcal{L}_{j}(\mathcal{J}(\mathcal{B}),\phi_{j})+\mathcal{L}_{i}(\mathcal{J}(\mathcal{B}),\phi_{i})\right)\right),$
$=\sum_{j\in\mathcal{B}}\left(\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right)\mathcal{L}_{j}(\mathcal{J}(\mathcal{B}),\phi_{j})+\sum_{i\in\mathcal{B}}\left(\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}\frac{1}{c_{ij}}\right)\mathcal{L}_{i}(\mathcal{J}(\mathcal{B}),\phi_{i}).$
Noting that $c_{ij}=c_{ji}$, by swapping the indices of the second term in the above summand, the second term is identical to the first term. Therefore, we deduce that
$\mathbb{E}_{\theta}[\mathcal{L}(\{\theta=\phi_{j}\},\mathcal{B})\;|\;j\in\mathcal{B}]\leq 2\sum_{j\in\mathcal{B}}\left(\left(\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right)\mathcal{L}_{j}(\mathcal{J}(\mathcal{B}),\phi_{j})\right).$
Now, using the fact that $\max_{j\in\mathcal{B}}\left(\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right)$ is a uniform upper bound for the multiplier in the above expression, we bring it outside the summation as a constant. Therefore, we deduce that
$\mathbb{E}_{\theta}[\mathcal{L}(\{\theta=\phi_{j}\},\mathcal{B})\;|\;j\in\mathcal{B}]\leq\left(\max_{j\in\mathcal{B}}\frac{2}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}\right)\mathcal{L}(\mathcal{J}(\mathcal{B}),\mathcal{B}),$ (5)
where we have used the fact that $\mathcal{L}(\mathcal{J}(\mathcal{B}),\mathcal{B})=\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\mathcal{J}(\mathcal{B}),\phi_{i})$. Since $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$, the covering $\mathcal{J(B)}$ is the loss minimizing service among all services in $\Theta_{\rm OPT}$ for all points in $\mathcal{B}$. Thus, we have that
$\mathbb{E}_{\theta}[\mathcal{L}(\{\theta=\phi_{j}\},\mathcal{B})\;|\;j\in\mathcal{B}]\leq\left(\max_{j\in\mathcal{B}}\frac{2}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}\right)\mathcal{L}(\Theta_{\rm OPT},\mathcal{B}).$
This concludes the proof. ∎
Lemma 2.
Let $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$ be an arbitrary cluster in $\mathcal{C}(\Theta_{\rm OPT})$, and let $\Theta_{t}\subset\mathbb{R}^{d}$ denote the parameters for a set of preexisting $t$ arbitrary services. Consider a new clustering $\Theta_{t+1}=\Theta_{t}\cup\theta$ where $\theta=\phi_{j}$ is a random service centered on user $j\in[n]$ selected with probability $P(\theta=\phi_{j})\propto\mathcal{L}_{j}(\Theta_{t},\phi_{j})$. Then, the expected loss of $\mathcal{B}$, conditioned on $j\in\mathcal{B}$, satisfies
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]\leq\frac{1}{\min_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}c_{ji}}\left(\max_{j\in\mathcal{B}}\frac{2}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}\right)\mathcal{L}(\Theta_{\rm OPT},\mathcal{B}).$
Proof.
Given that we are choosing a user from $\mathcal{B}$, the conditional probability that we center the new service on some fixed $\phi_{j}$ is precisely $\mathcal{L}(\Theta_{t},\phi_{j})/(\sum_{i\in\mathcal{B}}\mathcal{L}(\Theta_{t},\phi_{i}))$. After adding $\phi_{j}$ to the list of services, a user $i$ will have loss $\min\{\mathcal{L}_{i}(\Theta_{t},\phi_{i}),\mathcal{L}_{i}(\phi_{j},\phi_{i})\}$. Therefore we deduce that
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]$ $=\sum_{j\in\mathcal{B}}P(\theta=\phi_{j}|\;\theta\in\mathcal{B})\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\Theta_{t}\cup\theta,\phi_{i}),$ (6)
$=\sum_{j\in\mathcal{B}}\frac{\mathcal{L}_{j}(\Theta_{t},\phi_{j})}{\sum_{l\in\mathcal{B}}\mathcal{L}_{l}(\Theta_{t},\phi_{l})}\sum_{i\in\mathcal{B}}\min\{\mathcal{L}_{i}(\Theta_{t},\phi_{i}),\mathcal{L}_{i}(\phi_{j},\phi_{i})\}.$
By Assumption 2.$(i)$, for any $\theta\in\mathbb{R}^{d}$ we have that
$c_{ji}\mathcal{L}_{j}(\theta,\phi_{j})\leq\mathcal{L}_{i}(\theta,\phi_{i})+\mathcal{L}_{i}(\phi_{j},\phi_{i}).$
We utilize the fact that given two functions $f,g:\Theta\rightarrow\mathbb{R}$, if $f(\theta)\leq g(\theta)\;\forall\theta\in\Theta$, then $\min_{\theta\in\Theta}f(\theta)\leq\min_{\theta\in\Theta}g(\theta)$. Hence, the following implication holds:
$c_{ji}\min_{\theta\in\Theta}\mathcal{L}_{j}(\theta,\phi_{j})\leq\min_{\theta\in\Theta}\mathcal{L}_{i}(\theta,\phi_{i})+\mathcal{L}_{i}(\phi_{j},\phi_{i})\ \implies\ c_{ji}\mathcal{L}_{j}(\Theta_{t},\phi_{j})\leq\mathcal{L}_{i}(\Theta_{t},\phi_{i})+\mathcal{L}_{i}(\phi_{j},\phi_{i}).$
Summing over all $i\in\mathcal{B}$, we get that
$\mathcal{L}_{j}(\Theta_{t},\phi_{j})\leq\frac{1}{\sum_{i\in\mathcal{B}}c_{ji}}\left(\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\Theta_{t},\phi_{i})+\mathcal{L}_{i}(\phi_{j},\phi_{i})\right)$
Applying this to $\mathcal{L}_{j}(\Theta_{t},\phi_{j})$ on the right hand side of (6), we have that
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]$ $\leq\sum_{j\in\mathcal{B}}\frac{1}{\sum_{i\in\mathcal{B}}c_{ji}}\left(1+\frac{\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\phi_{j},\phi_{i})}{\sum_{l\in\mathcal{B}}\mathcal{L}_{l}(\Theta_{t},\phi_{l})}\right)\sum_{i\in\mathcal{B}}\min\{\mathcal{L}_{i}(\Theta_{t},\phi_{i}),\mathcal{L}_{i}(\phi_{j},\phi_{i})\}.$
Since $\min\{a,b\}\leq a$, we further deduce that
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]$ $\leq\sum_{j\in\mathcal{B}}\frac{1}{\sum_{i\in\mathcal{B}}c_{ji}}\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\phi_{j},\phi_{i})$
$\qquad+\sum_{j\in\mathcal{B}}\frac{1}{\sum_{i\in\mathcal{B}}c_{ji}}\frac{\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\phi_{j},\phi_{i})}{\sum_{l\in\mathcal{B}}\mathcal{L}_{l}(\Theta_{t},\phi_{l})}\sum_{i\in\mathcal{B}}\mathcal{L}(\Theta_{t},\phi_{i}),$
$=\sum_{j\in\mathcal{B}}\frac{2}{\sum_{i\in\mathcal{B}}c_{ij}}\sum_{i\in\mathcal{B}}\mathcal{L}_{i}(\phi_{j},\phi_{i}).$
Now, we use the upper bound $\frac{1}{\min_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}c_{ji}}$ for the multiplier to get
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]\leq\frac{2}{\frac{1}{|\mathcal{B}|}\min_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}c_{ji}}\sum_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\mathcal{L}_{i}(\phi_{j},\phi_{i}).$
Note that $\sum_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\mathcal{L}_{i}(\phi_{j},\phi_{i})$ is essentially the loss on choosing users from $\mathcal{B}$ uniformly randomly and centering the service on the chosen user. Plugging in the expression in (5) from the proof of Lemma 1, we have that
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]\leq\frac{4}{\min\limits_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum\limits_{i\in\mathcal{B}}c_{ji}}\left(\max_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right)\mathcal{L}(\mathcal{J}(\mathcal{B}),\mathcal{B}).$ (7)
Since $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$, the covering $\mathcal{J(B)}$ is the loss minimizing service among all available services in $\Theta_{\rm OPT}$ for all points in $\mathcal{B}$. Therefore, we deduce that
$\mathbb{E}_{\theta}[\mathcal{L}(\Theta_{t}\cup(\theta=\phi_{j}),\mathcal{B})\;|\;j\in\mathcal{B}]\leq\frac{4}{\min\limits_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum\limits_{i\in\mathcal{B}}c_{ji}}\left(\max_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right)\mathcal{L}(\Theta_{\rm OPT},\mathcal{B}),$
which concludes the proof. ∎
The following Lemma is an induction relating the losses on covered and uncovered clusters.
Lemma 3.
Let $\Theta_{t}\subset\mathbb{R}^{d}$ denote the parameters of a set of $t$ arbitrary services. Consider $u>0$ uncovered clusters from $\mathcal{C}(\Theta_{\rm OPT})$, denoted by $\mathcal{U}_{t}$, and let $\mathcal{H}_{t}$ denote the covered clusters. Suppose we add $v\leq u$ random services to $\Theta_{t}$, chosen with probability proportional to their current loss—i.e., $\mathcal{L}(\Theta_{t},\phi_{j})$—and let $\Theta_{t+v}$ denote the the resulting set of services. The following estimate holds:
$\mathbb{E}_{\Theta_{t+v}}[\mathcal{L}(\Theta_{t+v},[n])]\leq\left(\mathbb{E}_{\Theta_{t}}[\mathcal{L}\left(\Theta_{t},\mathcal{H}_{t}\right)]+K_{\rm OPT}\cdot\mathcal{L}\left(\Theta_{\rm OPT},\mathcal{U}_{t}\right)\right)\cdot\left(1+S_{v}\right)+\frac{u-v}{u}\cdot\mathbb{E}_{\Theta_{t}}[\mathcal{L}\left(\Theta_{t},\mathcal{U}_{t}\right)],$
where $K_{\rm OPT}$ is as defined in (2) and $S_{v}=\left(1+\frac{1}{2}+\ldots+\frac{1}{v}\right)$ is the harmonic series.
Proof.
Replacing the factor 8 in Lemma 3.3 from (Arthur and Vassilvitskii, 2007) by $K_{\rm OPT}$ gives us the desired result.
With the preceding technical lemmas, we are now ready to prove the main theorem which we restate for convenience. See 1
Proof.
Consider $t=1$, and $\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})$, the cluster in which the first chosen user belongs. Applying Lemma 3 with $v=u=k-1$ and the fact that $\mathcal{B}$ is the only covered cluster, we have that
$\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[n])]\leq\left(\mathbb{E}_{\Theta_{1}}[\mathcal{L}\left(\Theta_{1},\mathcal{B}\right)]+K_{\rm OPT}\cdot\mathcal{L}\left(\Theta_{\rm OPT},[n]-\mathcal{B}\right)\right)\cdot\left(1+S_{k-1}\right).$
Observe that
$\mathbb{E}_{\Theta_{1}}[\mathcal{L}\left(\Theta_{1},\mathcal{B}\right)]$ $\leq K_{\rm OPT}\cdot\mathcal{L}\left(\Theta_{\rm OPT},\mathcal{B}\right)],$
by Lemma 1. Moreover, we have that
$\mathcal{L}\left(\Theta_{\rm OPT},[n]-\mathcal{B}\right)=\mathcal{L}\left(\Theta_{\rm OPT},[n]\right)-\mathcal{L}\left(\Theta_{\rm OPT},\mathcal{B}\right).$
Using these two expression and noting $S_{k-1}\leq 1+\log k$, we get the stated result. ∎
Appendix C Fair Initialization
See 2
Proof.
For any $\Theta\subset\mathbb{R}^{d}$, we have that
$\Phi(\Theta,\mathcal{A})$ $=\max_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}\leq\frac{1}{\min_{i\in[m]}|\mathcal{A}_{i}|}\sum_{i\in[m]}\mathcal{L}(\Theta,\mathcal{A}_{i})=\frac{1}{\min_{i\in[m]}|\mathcal{A}_{i}|}\mathcal{L}(\Theta,[n]).$
This implies for $\Theta_{k}$ the output of Algorithm 1, we have that
$\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq\frac{1}{\min_{i\in[m]}|\mathcal{A}_{i}|}\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[n])].$
Hence, we have that
$\mathcal{L}(\Theta_{\rm fair},[n])$ $=\sum_{i\in[m]}\mathcal{L}(\Theta_{\rm fair},\mathcal{A}_{i})\leq\max_{i\in[m]}|\mathcal{A}_{i}|\sum_{i\in[m]}\max_{i\in[m]}\frac{\mathcal{L}(\Theta_{\rm fair},\mathcal{A}_{i})}{|\mathcal{A}_{i}|}=\max_{i\in[m]}|\mathcal{A}_{i}|m\cdot\Phi(\Theta_{\rm fair},\mathcal{A}).$
Now, since $\Theta_{\rm OPT}$ is the minimizer of $\mathcal{L}(\cdot,[n])$ and $\Theta_{\rm fair}$ the minimizer of $\Phi(\cdot,\mathcal{A})$, we have that
$\mathcal{L}(\Theta_{\rm OPT},[n])$ $\leq\max_{i\in[m]}|\mathcal{A}_{i}|m\cdot\Phi(\Theta_{\rm fair},\mathcal{A})$
$\leq\max_{i\in[m]}|\mathcal{A}_{i}|m\cdot\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]$
$\leq\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}m\cdot\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[n])].$
Using Theorem 1 for the guarantee on $\Theta_{k}$, we conclude the proof. ∎
C.1 What if User Demographic Identities are known?
We assume the scenario when the user demographic identities are known a priori. Let $\gamma:[n]\rightarrow\mathcal{A}$ be a mapping, which maps a user $i$ to its demographic group $\gamma(i)$, and let $|\gamma(i)|$ denote the size of the demographic group user $i$ belongs to.
We define the weighted total loss (where each users’ loss is divided by its group size) as:
$\mathcal{G}(\Theta,[n])$ $=\sum_{i\in[n]}\frac{\mathcal{L}_{i}(\Theta,\phi_{i})}{|\gamma(i)|}=\sum_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}.$
Note that this can also be interpreted as the sum of average losses across different demographic groups.
We introduce Algorithm 2 with a couple of changes from Algorithm 1.
  • Instead of uniformly picking a user at random, we are picking a user $i$ with probability proportional to $\frac{1}{|\gamma(i)|}$. This is equivalent to saying, the probability of the first user being in some group $\mathcal{A}_{j}\in\mathcal{A}$ is
    $\frac{\sum_{i\in\mathcal{A}_{j}}\frac{1}{|\gamma(i)|}}{\sum_{j\in[m]}\sum_{i\in\mathcal{A}_{j}}\frac{1}{|\gamma(i)|}}=\frac{1}{m}.$
    That is, the first user is equally likely to belong to one of the $m$ groups.
  • A user $i$ at any time step $t$ is selected with probability proportional to $\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})/|\gamma(i)|$. This implies that a user from some group $\mathcal{A}_{j}\in\mathcal{A}$ is likely to be picked with probability proportional to the average demographic loss—i.e.,
    $\sum_{i\in\mathcal{A}_{j}}\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})/|\gamma(i)|=\frac{\mathcal{L}(\Theta_{t-1},\mathcal{A}_{j})}{|\mathcal{A}_{j}|}.$
We first state the approximation ratio of Algorithm 2 on the objective $\mathcal{G}(\Theta,[n])$.
Lemma 4.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and unknown associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 1 and 2. Let $\Theta_{\rm scaled}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the total loss and $\mathcal{C}(\Theta_{\rm scaled})$ the resulting partitioning of users. If Algorithm 2 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
$\mathbb{E}_{\Theta_{k}}[\mathcal{G}(\Theta_{k},[n])]$ $\leq K_{\rm fair}(2+\log k)\cdot\mathcal{G}(\Theta_{\rm scaled},[n]),$
where the expectation is taken over the randomization of the algorithm and $K_{\rm fair}$ is equal to
$\max_{\mathcal{B}\in\mathcal{C}(\Theta_{\rm scaled})}\frac{4}{\min\limits_{j\in\mathcal{B}}\sum\limits_{i\in\mathcal{B}}\frac{c_{ij}}{|\gamma(i)|}}\left(\max_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}|\gamma(i)|}\right).$ (8)
The proof is similar as Theorem 1, with the introduction of $\frac{1}{|\gamma(i)|}$’s when deriving Lemma 1 and 2.
Algorithm 2 Fair AcQUIre- Fair Adaptively Querying Users for Initialization
1:  Input: Set of users $[n]$, number of services $k$, Demographic groups $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}$, Map users to demographic groups $\gamma:[n]\rightarrow\mathcal{A}$
2:  Choose a user $i$ uniformly randomly from $[n]$ with probability $\propto\frac{1}{|\gamma(i)|}$.
3:  Query user $i$’s preference $\phi_{i}$, set the first service $\Theta_{1}=\phi_{i}$.
4:  for $t\in\{2,\ldots,k\}$ do
5:     User behavior: Collect user losses on existing services $\Theta_{t-1}$ : $\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[n]}$.
6:     User Selection: Sample $l$ from $[n]$ with probability $P(l=i)\propto\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})/|\gamma(i)|$.
7:     New service: Query user $l$’s preference $\phi_{l}$.
8:     $\Theta_{t}=\Theta_{t-1}\cup\phi_{l}$.
9:  end for
10:  Return $\Theta_{k}$
Theorem 4.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and unknown associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 1 and 2. Suppose these users belong to $m$ demographic groups $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. Let $\Theta_{\rm fair}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the fairness objective $\Phi$ given in (3). If Algorithm 2 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
$\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq mK_{\rm fair}(2+\log k)\cdot\Phi(\Theta_{\rm fair},\mathcal{A}),$
where the expectation is taken over the randomization of the algorithm and $K_{\rm fair}$ is defined in (8).
Proof.
For any $\Theta\subset\mathbb{R}^{d}$, we have that
$\Phi(\Theta,\mathcal{A})$ $=\max_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}\leq\sum_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}=\mathcal{G}(\Theta,[n]).$
For $\Theta_{k}$—namely the output of Algorithm 1—the above expression implies that
$\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq\mathbb{E}_{\Theta_{k}}[\mathcal{G}(\Theta_{k},[n])].$
We also have that
$\mathcal{G}(\Theta_{\rm fair},[n])$ $=\sum_{i\in[m]}\frac{\mathcal{L}(\Theta_{\rm fair},\mathcal{A}_{i})}{|\mathcal{A}_{i}|}\leq\sum_{i\in[m]}\max_{i\in[m]}\frac{\mathcal{L}(\Theta_{\rm fair},\mathcal{A}_{i})}{|\mathcal{A}_{i}|}=m\cdot\Phi(\Theta_{\rm fair},\mathcal{A}).$
Let $\Theta_{\rm scaled}$ be the minimizer of $\mathcal{G}(\cdot,[n])$ and $\Theta_{\rm fair}$ be the minimizer of $\Phi(\cdot,\mathcal{A})$. Then, we have that
$\mathcal{G}(\Theta_{\rm scaled},[n])\leq\mathcal{G}(\Theta_{\rm fair},[n])\leq m\cdot\Phi(\Theta_{\rm fair},\mathcal{A})\leq m\cdot\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq m\cdot\mathbb{E}_{\Theta_{k}}[\mathcal{G}(\Theta_{k},[n])].$
Now using Lemma 4 for the guarantee on $\Theta_{k}$ for $\mathcal{G}(\cdot,[n])$, we conclude the proof. ∎
Remark 2.
If all $c_{ij}$’s are identically equal to some $c>0$, $K_{\rm fair}=\frac{4}{c^{2}}$ and the approximation ratio of Algorithm 2 for the fair objective is $4m(2+\log k)/c^{2}$. Meanwhile, Algorithm 1 would have an approximation ratio
$4m\cdot\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}\cdot\frac{(2+\log k)}{c^{2}}.$
Appendix D Proof of Theorem 3
Let the empirical loss—which is a finite sample average regression loss—be denoted
^$\widehat{\mathcal{L}}_{i}(\theta,\phi_{i})=\frac{1}{n_{i}}\sum_{j\in[n_{i}]}(\theta^{\top}x_{i}^{j}-y_{i}^{j})^{2},$
where $\{(x_{i}^{j},y_{i}^{j})\}_{j\in[n_{i}]}$ are private unknown features and scores of the users, respectively.
Lemma 5.
Under Assumptions 3 and 5, each subpopulation empirical loss ^$\widehat{\mathcal{L}}_{i}$ satisfies Assumption 1 and Assumption 2 with
$c_{ij}=\frac{1}{2}\lambda_{\mathrm{min}}\left(\frac{\sum_{l\in[n_{i}]}x_{i}^{l}(x_{i}^{l})^{\top}}{n_{i}},\frac{\sum_{l\in[n_{j}]}x_{j}^{l}(x_{j}^{l})^{\top}}{n_{j}}\right).$
Proof.
Fix a subpopulation index $i$. We start by showing ^$\widehat{\mathcal{L}}_{i}$ satisfies Assumption 1. We use the following result: if $p\leq d$ random vectors in $\mathbb{R}^{d}$ are independently drawn from a distribution that is absolutely continuous with respect to the Lebesgue measure, then they are almost surely linearly independent. By Assumption 5, applying this result to our scenario, if we draw $n_{i}\geq d$ features independently from a sub-Gaussian distribution to form a feature matrix $\mathbf{X}_{i}\in\mathbb{R}^{n_{i}\times d}$, then it is almost surely full column rank.
Therefore, we compute $\phi_{i}=\mathbf{X}_{i}^{\dagger}y_{i}$ since $\mathbf{X}_{i}^{\dagger}\mathbf{X}_{i}=\mathbf{I}_{p}$ when $\mathbf{X}_{i}$ is full column rank. The subpopulation $i$’s empirical loss is compactly written as ^$\widehat{\mathcal{L}}_{i}(\theta,\phi_{i})=\|\theta-\phi_{i}\|_{\mathbf{A}_{i}}^{2}$, where $\mathbf{A}_{i}=\frac{1}{n_{i}}\sum_{l\in[n_{i}]}x_{i}^{l}(x_{i}^{l})^{\top}$. Thus ^$\widehat{\mathcal{L}}_{i}(\theta,\phi_{i})$ satisfies Assumption 1.
We now show that ^$\widehat{\mathcal{L}}_{i}$ satisfies Assumption 2 with $c_{ij}=\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})$. To this end, we find the largest $c\in\mathbb{R}_{+}$ such that $c(u^{\top}\mathbf{A}_{i}u)\leq u^{\top}\mathbf{A}_{j}u$ and $c(u^{\top}\mathbf{A}_{j}u)\leq u^{\top}\mathbf{A}_{i}u$ for all $u\in\mathbb{R}^{d}$.
Rearranging the inequality, this problem is the same as finding the largest $c\in\mathbb{R}_{+}$ such that
$u^{\top}(\mathbf{A}_{i}-c\mathbf{A}_{j})u\geq 0\quad\text{and}\quad\mathbf{A}_{j}-c\mathbf{A}_{i})u\geq 0\quad\forall\ u\in\mathbb{R}^{d},$
or equivalently,
$\mathbf{A}_{i}-c\mathbf{A}_{j}\succeq 0\quad\text{and}\quad\mathbf{A}_{j}-c\mathbf{A}_{i}\succeq 0.$
Therefore, finding such a constant $c$ is easily reformulated as the following optimization problem:
$\max\{c~{}|\ \mathbf{A}_{i}-c\mathbf{A}_{j}\succeq 0,\mathbf{A}_{j}-c\mathbf{A}_{i}\succeq 0\}=\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j}).$
There are two key tools we use to finish the argument.
  • For any $u\in\mathbb{R}^{d}$, $\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})\|u\|_{\mathbf{A}_{i}}^{2}\leq\|u\|_{\mathbf{A}_{j}}^{2}$.
  • For any norm triangle inequality gives us $\|\theta-\phi_{i}\|\leq\|\theta-\phi_{j}\|+\|\phi_{i}-\phi_{j}\|$. Then we use power mean inequality, i.e. $a\leq b+c\implies a^{2}\leq 2(b^{2}+c^{2})$.
The following shows Assumption 2.$(i)$ holds with $c_{ij}=\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})$:
$\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})\widehat{\mathcal{L}}_{i}(\theta,\phi_{i})$ $=\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})\|\theta-\phi_{i}\|_{\mathbf{A}_{i}}^{2}$
$\leq\frac{1}{2}\|\theta-\phi_{i}\|_{\mathbf{A}_{j}}^{2}$
$\leq\|\theta-\phi_{j}\|^{2}_{\mathbf{A}_{j}}+\|\phi_{i}-\phi_{j}\|^{2}_{\mathbf{A}_{j}}$
=^$=\widehat{\mathcal{L}}_{j}(\theta,\phi_{j})+\widehat{\mathcal{L}}_{j}(\phi_{i},\phi_{j}).$
Analogously, the following shows Assumption 2.$(ii)$ holds with $c_{ij}=\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})$:
$\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})\widehat{\mathcal{L}}_{i}(\phi_{j},\phi_{i})$ $=\frac{1}{2}\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})\|\phi_{j}-\phi_{i}\|_{\mathbf{A}_{i}}^{2}$
$\leq\lambda_{\rm\min}(\mathbf{A}_{i},\mathbf{A}_{j})\left(\|\theta-\phi_{j}\|_{\mathbf{A}_{i}}^{2}+\|\theta-\phi_{i}\|_{\mathbf{A}_{i}}^{2}\right)$
$\leq\|\theta-\phi_{j}\|_{\mathbf{A}_{j}}^{2}+\|\theta-\phi_{i}\|^{2}_{\mathbf{A}_{i}}$
$=\mathcal{L}_{j}(\theta,\phi_{j})+\mathcal{L}_{i}(\theta,\phi_{i}).$
This concludes the proof. ∎
With the preceding technical lemma in place, we now are ready to prove Theorem 3. See 3
Proof.
Lemma 5 gives us that the subpopulation $i$’s empirical loss for a regressor $\theta\in\mathbb{R}^{d}$ can be written as $\|\theta-\phi_{i}\|^{2}_{\mathbf{A}_{i}}$. This is a random quantity since $\mathbf{A}_{i}$ is a sample average covariance of randomly chosen features. We analyse this term next.
For a random feature $x\in\mathbb{R}^{d}$ in subpopulation $i$, given a fixed center $\theta\in\mathbb{R}^{d}$, the term $\frac{(\theta-\phi_{i})^{\top}x}{\|\theta-\phi_{i}\|_{\Sigma_{i}}}$ is sub-Gaussian with variance proxy $\sigma_{i}^{2}$ by Assumption 3. The square of a sub-Gaussian random variable, $\frac{(\theta-\phi_{i})^{\top}xx^{\top}(\theta-\phi_{i})}{\|\theta-\phi_{i}\|^{2}_{\Sigma_{i}}}$ is sub-exponential [^7] A random variable $x$ is subexponential with parameters $(v^{2},\alpha)$ if $\mathbb{E}[\exp(\lambda x)]\leq\exp(\frac{\nu^{2}\lambda^{2}}{2})\;\forall|\lambda|\leq\frac{1}{\alpha}$. with parameters $\nu_{i}^{2}=32\sigma_{i}^{4},\alpha_{i}=4\sigma_{i}^{2}$ (see (Honorio and Jaakkola, 2014, Appendix B)).
Since subpopulation $i$’s empirical loss is an average of $n_{i}$ samples, the empirical loss $\|\theta-\phi_{i}\|^{2}_{\mathbf{A}_{i}}$ is a subexponential variable with parameters $\frac{\nu_{i}^{2}}{n_{i}}\|\theta-\phi_{i}\|^{4}_{\Sigma_{i}},\frac{\alpha_{i}}{\sqrt{n_{i}}}\|\theta-\phi_{i}\|^{2}_{\Sigma_{i}}$. This is because the variance of sample average scales as $\frac{1}{\sqrt{n_{i}}}$, and the constant $\|\theta-\phi_{i}\|^{2}_{\Sigma_{i}}$ scales the subexponential parameters appropriately.
Now, given a set of $k$ arbitrary services $\Theta=\{\theta_{1},\ldots,\theta_{k}\}\subset\mathbb{R}^{d}$, let $\theta_{\gamma(i)}\in\Theta$ be an arbitrary service chosen by subpopulation $i$. The sum of empirical losses on based on such choices is also subexponential with parameters $\nu^{2}=\sum_{i\in[N]}\frac{\nu_{i}^{2}}{n_{i}}\|\theta-\phi_{i}\|^{4}_{\Sigma_{i}}$ and $\alpha=\max_{i\in[N]}\frac{\alpha_{i}}{\sqrt{n_{i}}}\|\theta-\phi_{i}\|^{2}_{\Sigma_{i}}$. Using Chernoff bound (Vershynin, 2018), we have that
$P\left\{|\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\mathbf{A}_{i}}-\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}|\right\}\geq t]\leq 2\exp\left(-\frac{t^{2}}{2\nu^{2}}\right)\quad\text{where}\ 0\leq t\leq\frac{\nu^{2}}{\alpha}.$ (9)
Note that the probability bound is is increasing in $\nu$, we thus take an upper bound on $\nu^{2}$ to upper bound this probability:
$\nu^{2}$ $=\sum_{i\in[N]}\frac{\nu_{i}^{2}}{n_{i}}\|\theta-\phi_{i}\|^{4}_{\Sigma_{i}}=32\sum_{i\in[N]}\frac{\sigma_{i}^{4}}{n_{i}}\|\theta_{\gamma(i)}-\phi_{i}\|^{4}_{\Sigma_{i}}\leq 32\left(\sum_{i\in[N]}\frac{\sigma_{i}^{8}}{n_{i}^{2}}\right)^{\frac{1}{2}}\left(\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{8}_{\Sigma_{i}}\right)^{\frac{1}{2}},$
where the last inequality follows from Cauchy-Schwarz. Further, noticing that $\sum_{i\in n}a_{i}^{4}\leq(\sum_{i\in n}a_{i}^{2})^{2}$, and applying it for the last term twice, we deduce that
$\nu^{2}\leq 32\left(\sum_{i\in[N]}\frac{\sigma_{i}^{8}}{n_{i}^{2}}\right)^{\frac{1}{2}}\left(\sum_{i\in[n]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}\right)^{2}.$
Choosing $t=\epsilon\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}$ in (9), where $0\leq\epsilon\leq\mathcal{O}(\min_{i\in[N]}\frac{\sigma_{i}^{2}}{\sqrt{n_{i}}})$, we have that
$P\left[|\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\mathbf{A}_{i}}-\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}|\geq\epsilon\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}\right]$
$\qquad\geq 2\exp{\left(-\frac{\epsilon^{2}}{64}\left({\sum_{i\in[N]}\frac{\sigma_{i}^{8}}{n_{i}^{2}}}\right)^{-1/2}\right)}.$
With high probability, at least $1-2\exp{\left(-\frac{\epsilon^{2}}{64}\left({\sum_{i\in[N]}\frac{\sigma_{i}^{8}}{n_{i}^{2}}}\right)^{-1/2}\right)}$, for all $\Theta\subset\mathbb{R}^{d}$, we have that
$(1-\epsilon)\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}\leq\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\mathbf{A}_{i}}\leq(1+\epsilon)\sum_{i\in[N]}\|\theta_{\gamma(i)}-\phi_{i}\|^{2}_{\Sigma_{i}}.$
The rest of the proof is deterministic, thus it is assumed everything follows with probability at least $1-2\exp{\left(-\frac{\epsilon^{2}}{64}\left({\sum_{i\in[N]}\frac{\sigma_{i}^{8}}{n_{i}^{2}}}\right)^{-1/2}\right)}$. Define $\Theta^{N}=\Theta\times\cdots\times\Theta$—i.e., Cartesian product of the set $\Theta$ $N$-times. Given two functions $f,g:\Theta^{N}\rightarrow\mathbb{R}$, if $f(\theta)\leq g(\theta)$ for all $\theta\in\Theta^{N}$, then $\min_{\theta\in\Theta^{N}}f(\theta)\leq\min_{\theta\in\Theta^{N}}g(\theta)$. Therefore, we deduce that
$(1-\epsilon)\sum_{i\in[N]}\min_{j\in[k]}\|\theta_{j}-\phi_{i}\|^{2}_{\Sigma_{i}}$ $\leq\sum_{i\in[N]}\min_{j\in[k]}\|\theta_{j}-\phi_{i}\|^{2}_{\mathbf{A}_{i}}\leq(1+\epsilon)\sum_{i\in[N]}\min_{j\in[k]}\|\theta_{j}-\phi_{i}\|^{2}_{\Sigma_{i}}.$ (10)
Hence, given the same set of services $\Theta$, the total empirical loss can be bounded with respect to the total expected loss as follows:
$(1-\epsilon)\mathcal{L}(\Theta,[N])\leq\widehat{\mathcal{L}}(\Theta,[N])\leq(1+\epsilon)\mathcal{L}(\Theta,[N]).$
Recall $\Theta_{\rm OPT}$ denotes the optimal solution for the total expected loss and let $\widehat{\Theta}_{\rm OPT}$ denote the optimal solution for the total empirical loss. Let $\Theta_{k}$ be the output of Algorithm 1 on finite samples per subpopulations. Using the first inequality of (10), we have that
$(1-\epsilon)\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[N])]\leq\mathbb{E}_{\Theta_{k}}[\widehat{\mathcal{L}}(\Theta_{k},[N])].$ (11)
Also, noting that $\widehat{\Theta}_{\rm OPT}$ is the minimizer for ^$\widehat{\mathcal{L}}(\Theta,[N])$ and using the second inequality of (10), we get that
^$\widehat{\mathcal{L}}(\widehat{\Theta}_{\rm OPT},[N])\leq\widehat{\mathcal{L}}(\Theta_{\rm OPT},\Phi)\leq(1+\epsilon)\mathcal{L}(\Theta_{\rm OPT},\Phi).$ (12)
Now combining (11) and (12) with Theorem 1, we get the desired result—i.e.,
$\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[N])]\leq\frac{1+\epsilon}{1-\epsilon}K_{\rm OPT}(2+\log k)\cdot\mathcal{L}(\Theta_{\rm OPT},[N]).$
Setting $\delta=1-2\exp{\left(-\frac{\epsilon^{2}}{64}\left({\sum_{i\in[N]}\frac{\sigma_{i}^{8}}{n_{i}^{2}}}\right)^{-1/2}\right)}$, we get that $n_{i}=\Omega(\frac{\sigma_{i}^{4}\sqrt{N}\log{(2/\delta)}}{\epsilon^{2}})$. This concludes the proof.
Figure 4: Fair objective improvement for AcQUIre over the baseline across different demographics. We observe that there is atleast 15% improvement across sex demographics for a wide range of number of services. For racial demographics the improvement is 7-26%.
Figure 5: Average losses across different demographic groups for Fair AcQUIre (left,middle). Percentage improvement over baseline (right). We observe that Fair AcQUIre reduces disparity across different groups compared to the baseline.
Figure 6: Average Total Loss across different baselines. We observe that while the greedy algorithm does when the number of services is small, its performance starts degrading as the number of services grows. This is beacuse the greedy algorithm is susceptible to pick outliers. AcQUIre beats both the baselines across a large range of number of services.
Appendix E Experiment Details
Census Dataset.
The categorical features (schooling, marital status, migration status, citizenship) were first converted to one hot vectors and the continuous features (age, income) were scaled appropriately for better conditioning. We performed singular value decomposition on the features and retained the top ten components. The scores were taken as the log transform of the daily commute time in minutes. To form subpopulations, we split users based on their Public Use Microdata Area codes (zip code) and ensure each subpopulation belongs to one of the demographic groups considered. The percentage improvements of our algorithm are plotted in Figure 4.
We also run Fair AcQUIre (Algorithm 2) on the Census experiment where the demographic identities are known to the provider. As a baseline we used a weighted sampling of users where the weights are inversely proportional to the demographic sizes. This baseline ensures the expected number of users sampled per demographic is equal. The results are reported in Figure 5.
Additional Baseline.
We consider an additional baseline for initialization process. The greedy algorithm at every iteration picks the user with the largest loss, queries the picked user’s preference and sets the new service to this preference. We find that while the algorithm performs well when the number of services $k$ is small, however as $k$ increases the uniform sampling starts performing better than greedy (see Figure 6). This is because a greedy algorithm is susceptible to pick outliers, particularly when $k\ll n$ and these outliers don’t generalize well to the average population.
Movie Recommendation Dataset.
We use Surprise (a Python toolkit (Hug, 2020)) to perform our experiments. We split the total 10 million ratings into top (a) 200 movies, and (b) all other movies. We use the inbuilt nonnegative matrix factorization function of Surprise on (b) to get user and item embeddings. We cluster the users into 1000 subpopulations by running $k$-means on the obtained user embeddings we get. We evaluate algorithms for this experiment on the held out set (a) of the top 200 movies.